OpenStudy (anonymous):

If 10.1 of Fe are added to a solution of Copper (Cu) sulfate, how many grams of solid copper would form? 2Fe + 3CuSO4 -> FFe2(SO4)3 + 3Cu

OpenStudy (zpupster):

\[10.1g Fe( \frac{ 1 Mole Fe }{ 55.85 g Fe })(\frac{ 3 moles Cu }{ 2 moles Fe })(\frac{ 63.55 g Cu }{ 1mole Cu })\] how many moles Fe? then we see 2 moles Fe makes 3 moles Cu finally how many grams of Cu is that?

OpenStudy (anonymous):

\[\frac{ 1925.565 }{ 111.7 }\] \[= 17.238~\]

OpenStudy (zpupster):

good job!!

OpenStudy (anonymous):

thanks C: